A radioactive nucleus of Mass M emits a photon of frequency and the nucleus recoils. The recoil will be


1
Expert's answer
2021-05-04T11:06:19-0400

Energy of emitted photon (E)=hv

Photon it's momentum is given as "=\frac{E}{velocity of length}"

Which is not nothing but

"\frac{E}{c}or \frac{hv}{c}"

According to momentum conservation value of the momentum of recoiled nucleous should also be "\frac{hv}{c}" But in opposite directions to make net momentum zero before and after emission

Recoil energy [E2]="\frac{P^2}{2m}"

P is recoil momentum putting

"P=\frac{hv}{c}"

We get recoil energy as

"E=\frac{h^2v^2}{2mc^2}"

Where "v\rightarrow frequency"